15 is 3/5 of n

What is 3/5 of 15?

Answer by checkley77(12844)

15 is 3/5 of n
15=3/5*n
15=3n/5 cross multiply
3n=5*15
3n=75
n=75/3
n=25 ans.
Proof:
15=3/5*25
15=75/5
15=15
What is 3/5 of 15?
3/5*15=45/5=9 ans.